OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN I DONT WANT AN ANSWER. SCROLL DOWN ALL THE WAY, I STILL NEED HELP Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. Part B: How many liters of lemon juice are present in the bottle? Show your work.

OpenStudy (welshfella):

let x be the total volume of the liquid in the bottle so the volume of lemon juice = 0.60 x litres there is 1 liter of water so we have x = 0.60x + 1 so 0.40x = 1

OpenStudy (welshfella):

thats part A for part B 0.40 x = 1 x = 1/0.40 = 2.5 lires so there are 2.5 litres of liquid of which 2.5 - 1 litres are lemon juice there is 1.5 litres of lemon juice

OpenStudy (princeharryyy):

u still need help ?